Find the seasonalized forecast for Q2 of 2017.The sales trend has been modeled as: Sales = 255.00+ 2 * t , where t= time in quarters, beginning in Q1 2014. Seasonality for the four quarterly periods is given in the table below.

QUARTER SEASONAL FACTOR
Q #1 1.2
Q #2 1.18
Q #3 0.5
Q #4

** Round to 1 Decimal Place! **
Answer #1

Pgo Answers seasonal under for Quarter 4 =4-(1.2+1.18+0.5) =4-2.88 = 1.12 sales =25stat Deseasonalized forecast for Q2 for 20